# ConGeo: Robust Cross-view Geo-localization across Ground View Variations

### v1

@@ Acknowledgments
We investigate the effect of loss weights on the retrieval performance. In Table S4, model performance with different loss weights on North-aligned setting and ${}^{\circ}$ are reported. Results suggest that different loss weights may lead to slight performance waves. Specifically, when we gradually increase the $w_{1}$, $w_{2}$, and $w_{3}$, the performance of R@1 under the North-aligned setting will slightly decrease within 1 point, meanwhile, the limited FoV performance will be improved. When the single-modal loss weight $w_{1}$ and $w_{2}$ value are set higher, there is a significant improvement on FoV=$90^{\circ}$ with around 4 points on R@1.
@@
In the main paper, we demonstrate that ConGeo, as a learning objective, can be plugged into different base models, including the CNN-based model [5] and ViT-based model [4]. To better compare those two mainstream architectures for cross-view geo-localization, we use Sample4Geo as the base model and switch the backbone between CNN and ViT in Table S6 to test the corresponding performance on the North-aligned settings and ${}^{\circ}$. In other words, the only difference between these two base models is the encoder architecture: Sample4Geo-CNN is with ConvNeXt [15] and Sample4Geo-ViT is with Swin [37]. From the results, we can see that ConGeo brings considerable improvement to both architectures, between them, the improvement with the CNN-based backbone is larger. We also compare the effect of training FoV angle ($\alpha$) on different architectures. The results indicate that CNN-based architecture yields the best overall performance when setting $\alpha$ to ${}^{\circ}$. The ViT-based architecture, although shows less competitive performance compared to the CNN-based one, seems to be more robust with different training angles $\alpha$. Here, we present preliminary findings regarding the analysis of different architectures under unknown orientations and limited FoV settings. As outlined in the Limitations section (Section 7) of the main paper, a more comprehensive investigation into this direction could serve as an intriguing topic for future research.
Method North-aligned ${}^{\circ}$
@@
Method North-aligned ${}^{\circ}$
@@
Sample4Geo w/o Polar 98.7 99.8 99.9 3.5 9.8 26.7
ConGeo w/o Polar 98.3 99.5 99.9 55.9 79.0 90.9
@@
As we reviewed in the Related Works section (Section 2) in the main paper, polar transformation is commonly used to improve the models’ performance on the North-aligned setting [2, 4]. Polar transformation transfers the aerial images to the polar coordinates and narrows the view gap between aerial images and ground images. We compare the performance between the base model and ConGeo with the original satellite images or with the polar transformed satellite images in Table S6. Results indicate ConGeo can consistently improve the robustness, whether the aerial images are polar-transformed or not. Note that polar transformation enhances the spatial correspondences in the model. Therefore, compared with ConGeo without polar transformation, ConGeo with polar transformation is more likely to maintain strong performance in the North-aligned setting, while being less competitive in the limited FoV settings.
@@
The transfer analysis between CVUSA and CVACT datasets is often reported to show the generalization ability of the model on different datasets. We also report the transfer results of the base model and ConGeo on the unknown orientation setting and limited FoV setting in Table S7. Results suggest ConGeo can significantly improve the robustness of the base model (R@1% gains on average 20.6% on ${}^{\circ}$).
S6 Supplementary Results
@@
$\mathcal{L}_{\text{single-r}}$ $\mathcal{L}_{\text{single-q}}$ $\mathcal{L}_{\text{cross}}$ ${}^{\circ}$ ${}^{\circ}$ ${}^{\circ}$ ${}^{\circ}$
@@
Table S9: The full table of ablation studies on ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$, and ${}^{\circ}$ on the CVUSA dataset. “Shift” and “FoV” mean cyclic shift and FoV cropping. “Single” and “Cross” denote single-modal contrastive objective cross-modal contrastive objective, respectively.
Augmentation ${}^{\circ}$ ${}^{\circ}$ ${}^{\circ}$ ${}^{\circ}$
@@
Table S10: The full table of comparison with different data augmentation methods on ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$, and ${}^{\circ}$ on the CVUSA dataset. “Shift” applies random cyclical shift, “FoV” means applying random FoV cropping (from ${}^{\circ}$ to ${}^{\circ}$) to ground images as data augmentation for training, ‘Rotate” randomly rotating aerial images with an angle in {${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$} as data augmentation.
@@
As shown in Table S9, through the full results provided, we can further confirm the individual effectiveness of each component in the proposed ConGeo. Moreover, the combination of single-modal and cross-modal contrastive objectives can further boost the model performance indicating that they can complement each other’s alignment. Besides, we can see from the results that merely applying “Shift” transformations to the ground view images in single-modal and cross-modal contrastive objectives achieves the best performance when FoV=$360^{\circ}$, showing that the model performance on specific FoVs can be further improved by using the evaluation FoV for training like previous methods do.
@@
Our method may fail when very limited information is provided in the query image. The first case is that the query image is of a very small FoV. As can be seen from Table 1 of the main paper, the retrieval performance under very limited FoV (e.g. ${}^{\circ}$) is significantly lower than other settings. The second case is that the query image contains merely single and common layouts. As we mentioned in the Limitations Section in the main paper, some samples exhibit greater sensitivity to the robustness test, making it more challenging for the model to uphold its robustness on such samples. The model will struggle to distinguish the corresponding aerial image from similar counterparts given common layouts.
We show the examples on the right part of Fig. S2 when the rank of ground truth annotations are 1, 4000, and 8000. In general, when the query image only contains less discriminative grass or trees, the similarity score of the ground truth aerial image may be drowned in other similar candidates. Analysing the sensitivity of different samples can also help improve model robustness and can be an interesting direction for future research.
S8 Discussion
In this work, we propose ConGeo to boost the base models’ robustness across ground view variations. Without specialized training, ConGeo outperforms state-of-the-art methods on unknown orientation and limited FoV benchmarks and demonstrates adaptability to three different model base models and generalizability to various unseen ground view variations. However, as mentioned in the Limitations section (Section 7) of the main paper, several aspects and challenges remain to be addressed for future research.
First, compared to models trained specifically for the North-aligned setting, ConGeo faces an almost unavoidable slight performance drop in this setting. Indeed, when performing orientation-specific training, North-alignment is a key assumption and models often take shortcuts by using the spatial correspondence in the data stemming from this alignment. However, ConGeo aims to relax it and achieve robustness in more challenging settings, where such shortcuts will reduce the model’s generality. Moreover, it is also worth noting that the North-aligned setting is a special case of ${}^{\circ}$, where the orientation angle is set to ${}^{\circ}$. Therefore, the performance drop in the North-aligned setting is almost unavoidable when improving the unknown orientation setting, which also indicates a trade-off between orientation-specific or FoV-specific spatial correspondence and the robust semantic information across ground view variations. Despite this slight decrement, ConGeo serves as a practical solution to real-world scenarios: When the orientation is known, ConGeo performs comparably to the state-of-the-art methods, while it significantly improves the performance when it is unknown.
@@
Finally, in this work, we mainly focus on ground view orientation and reduced FoV variations. However, we also validate the effectiveness of the model under several unseen variations (Table 8 in the main paper), such as blur and zooming. Expanding the robust geo-localization across a broader range of ground view variations (e.g., nighttime or extreme weather conditions) can further widen the applicability of research to real-world use cases.
@@
Figure S2: Distribution of the rank of ground truth reference images in the retrieval results when FoV=$90^{\circ}$ (left) and some examples of query images according to the retrieval results of ConGeo.
Figure S3: More examples of the Grad-CAM activation maps of the base model and ConGeo on the North-aligned setting (the first row for each sample) and the unknown orientation setting (the second row for each sample). The orange box indicates the same area in different ground view variants.
Figure S4: More examples of the Grad-CAM activation maps of the base model (top row) and ConGeo (bottom row) on the North-aligned setting (a) and limited FoV setting (b) on the left. Corresponding patches from the two settings are shown on the right.

### v2

@@ Acknowledgments
We investigate the effect of loss weights on the retrieval performance. In Table S4, model performance with different loss weights on North-aligned setting and FoV=90 are reported. Results suggest that different loss weights may lead to slight performance waves. Specifically, when we gradually increase the $w_{1}$, $w_{2}$, and $w_{3}$, the performance of R@1 under the North-aligned setting will slightly decrease within 1 point, meanwhile, the limited FoV performance will be improved. When the single-view loss weight $w_{1}$ and $w_{2}$ value are set higher, there is a significant improvement on FoV=$90^{\circ}$ with around 4 points on R@1.
@@
In the main paper, we demonstrate that ConGeo, as a learning objective, can be plugged into different base models, including the CNN-based model [5] and ViT-based model [42]. To better compare those two mainstream architectures for cross-view geo-localization, we use Sample4Geo as the base model and switch the backbone between CNN and ViT in Table S6 to test the corresponding performance on the North-aligned settings and FoV=90. In other words, the only difference between these two base models is the encoder architecture: Sample4Geo-CNN is with ConvNeXt [19] and Sample4Geo-ViT is with Swin Transformer [18]. From the results, we can see that ConGeo brings considerable improvement to both architectures, between them, the improvement with the CNN-based backbone is larger. We also compare the effect of training FoV angle ($\alpha$) on different architectures. The results indicate that CNN-based architecture yields the best overall performance when setting $\alpha$ to 180. The ViT-based architecture, although shows less competitive performance compared to the CNN-based one, seems to be more robust with different training angles $\alpha$. Here, we present preliminary findings regarding the analysis of different architectures under unknown orientations and limited FoV settings. A more comprehensive investigation into this direction could serve as an intriguing topic for future research.
Method North-aligned FoV=90
@@
Method North-aligned FoV=90
@@
Sample4Geo w/o Polar 98.7 99.8 99.9 2.5 9.8 26.7
ConGeo w/o Polar 98.3 99.7 99.9 55.9 79.0 90.9
@@
As we reviewed in Section 2 in the main paper, polar transformation is commonly used to improve the models’ performance on the North-aligned setting [27, 42]. Polar transformation transfers the aerial images to the polar coordinates and narrows the view gap between aerial images and ground images. We compare the performance between the base model and ConGeo with the original satellite images or with the polar transformed satellite images in Table S6. Results indicate ConGeo can consistently improve the robustness, whether the aerial images are polar-transformed or not. Note that polar transformation enhances the spatial correspondences in the model. Therefore, compared with ConGeo without polar transformation, ConGeo with polar transformation is more likely to maintain strong performance in the North-aligned setting, while being less competitive in the limited FoV settings.
@@
The transfer analysis between CVUSA and CVACT datasets is often reported to show the generalization ability of the model on different datasets. We also report the transfer results of the base model and ConGeo on the unknown orientation setting and limited FoV setting in Table S7. Results suggest ConGeo can significantly improve the robustness of the base model (R@1% gains on average 20.6% on FoV=360 over the two settings).
S5.4 Different Paradigms of Aligning Modalities
ConGeo requires the alignment of different modalities. Besides contrastive learning (CL), other methods, e.g., direct feature alignment (FA), and redundancy reduction (RR), can also be used to align modalities. To investigate this, we also compare different paradigms of alignment modalities using InfoNCE [22] (CL), Barlow Twins [38] (RR) and Consine similarity (FA) to represent each paradigm respectively. The results shown in Table S8 suggest that aligning modalities helps to improve robustness, but due to the huge cross-modal view gap and single-view information asymmetry, FA fails to reduce the feature distance, RR struggles to find the shared features, while CL excels in learning by comparing.
Paradigm Method North-Aligned FoV=90
R@1 R@1% R@1 R@1%
FA Cosine Similarity 86.6 99.6 31.2 81.9
RR Barlow Twins [38] 85.1 99.2 41.4 82.1
CL [ConGeo] InfoNCE [22] 98.3 99.9 55.9 90.9
Table S8: Comparison of different paradigms of aligning modalities. “FA”, “RR” and “CL” denote direct alignment, redundancy reduction and contrastive learning, respectively.
Appendix S6 Supplementary Results
@@
$\mathcal{L}_{\text{single-r}}$ $\mathcal{L}_{\text{single-q}}$ $\mathcal{L}_{\text{cross}}$ FoV=360 FoV=180 FoV=90 FoV=70
@@
Table S10: The full table of ablation studies on FoV=70, 90, 180, and 360 on the CVUSA dataset. “Shift” and “FoV” mean cyclic shift and FoV cropping. “Single” and “Cross” denote single-view contrastive objective cross-view contrastive objective, respectively.
Augmentation FoV=360 FoV=180 FoV=90 FoV=70
@@
Table S11: The full table of comparison with different data augmentation methods on FoV=70, 90, 180, and 360 on the CVUSA dataset. “Shift” applies random cyclical shift, “FoV” means applying random FoV cropping (from 70 to 360) to ground images as data augmentation for training, ‘Rotate” randomly rotating aerial images with an angle in {90, 180, 270} as data augmentation.
@@
As shown in Table S10, through the full results provided, we can further confirm the individual effectiveness of each component in the proposed ConGeo. Moreover, the combination of single-view and cross-view contrastive objectives can further boost the model performance indicating that they can complement each other’s alignment. Besides, we can see from the results that merely applying “Shift” transformations (FoV=$360^{\circ}$) to the ground view images in single-view and cross-view contrastive objectives achieves the best performance, showing that the model performance on specific FoVs can be further improved by using the evaluation FoV for training like in previous methods.
@@
Our method may fail when very limited information is provided in the query image. The first case is that the query image is of a very small FoV. As can be seen from Table 1 of the main paper, the retrieval performance under very limited FoV (e.g. 70) is significantly lower than other settings. The second case is that the query image contains merely single and common layouts. Some samples exhibit greater sensitivity to the robustness test, making it more challenging for the model to uphold its robustness on such samples. The model will struggle to distinguish the corresponding aerial image from similar counterparts given common layouts.
We show the examples on the right part of Fig. S2 when the rank of ground truth annotations are 14,000, and 8,000. In general, when the query image only contains grass or trees, the similarity score of the ground truth aerial image may be drowned in other similar candidates. Analysing the sensitivity of different samples can also help improve model robustness and can be an interesting direction for future research.
Appendix S8 Discussion
In this work, we propose ConGeo to boost the base models’ robustness across ground view variations. Without specialized training, ConGeo outperforms state-of-the-art methods on unknown orientation and limited FoV benchmarks and demonstrates adaptability to three different base models and generalizability to various unseen ground view variations. However, as mentioned in Section 7 of the main paper, several aspects and challenges remain to be addressed for future research.
First, compared to models trained specifically for the North-aligned setting, ConGeo faces an almost unavoidable slight performance drop in this setting. Indeed, when performing orientation-specific training, North-alignment is a key assumption and models often take shortcuts by using the spatial correspondence in the data stemming from this alignment. However, ConGeo aims to relax it and achieve robustness in more challenging settings, where such shortcuts will reduce the model’s generality. Moreover, it is also worth noting that the North-aligned setting is a special case of FoV=360, where the orientation angle is set to 0. Therefore, the performance drop in the North-aligned setting is almost unavoidable when improving the unknown orientation setting, which also indicates a trade-off between orientation-specific or FoV-specific spatial correspondence and the robust semantic information across ground view variations. Despite this slight decrement, ConGeo serves as a practical solution to real-world scenarios: When the orientation is known, ConGeo performs comparably to the state-of-the-art methods, while the model significantly improves the performance when the orientation is unknown.
@@
Finally, in this work, we mainly focus on ground view orientation and reduced FoV variations. However, we also validate the effectiveness of the model under several unseen variations (Table 8 in the main paper), such as blur and zooming. Expanding the robust geo-localization across a broader range of ground view variations (e.g., nighttime or extreme weather conditions) can further widen the applicability of this research to real-world use cases.
@@
Figure S2: Distribution of the rank of ground truth reference images in the retrieval results when FoV=$90^{\circ}$ (left) and some examples of query images according to the retrieval results of ConGeo (right).
Figure S3: Additional examples of the Grad-CAM activation maps of the base model and ConGeo on the North-aligned setting (the first row for each sample) and the unknown orientation setting (the second row for each sample). The orange box indicates the same area in different ground view variants.
Figure S4: Additional examples of the Grad-CAM activation maps of the base model (top row) and ConGeo (bottom row) on the North-aligned setting (a) and limited FoV setting (b) on the left. Corresponding patches from the two settings are shown on the right.